2. What is the molarity of NaOH solution if 35.75 mL of this solution is required to neutralize 1.070 g KHP.

3. What is the molarity of HCl solution if 25.0 mL of this solution required 15.0 mL of 0.050 M NaOH for neutralization?


1
Expert's answer
2020-12-01T09:04:43-0500

2. HKC8H4O4 + NaOH → NaKC8H4O4 + H2O

M(HKC8H4O4) = 204.23 g/mol

n(HKC8H4O4) "= \frac{1.070}{204.23} = 0.00524 \; mol"

n(HKC8H4O4) = n(NaOH) = 0.00524 mol

Proportion:

35.75 mL – 0.00524 mol

1000 mL – x

"x = \frac{1000 \times 0.00524}{35.75} = 0.146 \; M"

Answer: 0.146 M NaOH


3. HCl + NaOH → NaCl + H2O

Proportion 1 (NaOH):

0.05 mol – 1000 mL

n mol – 15.0 mL

n(NaOH) = "\frac{0.05 \times 15.0}{1000} = 0.00075 \;mol"

n(HCl) = n(NaOH) = 0.00075 mol

Proportion 2 (HCl):

25.0 mL – 0.00075 mol

1000 mL – x

"x = \frac{1000 \times 0.00075}{25.0} = 0.03 \;M"

Answer: 0.03 M HCl